Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 18:32:03 on September 17, 2009
Entry number 290172
keyword=Run Plan for Tonight
Run Plan
- The plan for the night: Take production data! Watch the lumi widths
to see if anything changes. Please check the histograms on the online
data monitor with the reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Take Compton data! You will need to start a new Compton run every
two hours. The CODA screen to start and stop the Compton runs is in the
back room. The optimal background rate seems to be 100 Hz/muA, but if it
over 300 Hz/muA, call MCC and have them adjust the beam tune.
- If the background rates get too high, the Compton HV will trip off
and you won't be able to take data. If you cannot keep it going, page
Megan Friend (584-0038), and if you cannot get her, page Alexandre
Camsonne (584-5064).
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
- Change the half-wave plate state when we get to 1M good pairs.
Follow the instructions for doing this and running the flipper script
here: http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#ChangingHalfWavePlate
- If you lose beam for an extended period, do our beam
diagnostics to make sure that everything is good (the optimal values for
the raster, beam position, beam current, lumi RMS, and detector asymmetry
RMS are on the whiteboard).
- With an empty target, do harp scans, which are done by MCC on
request.
- Still with an empty target, do spot++ at 40 muA. Change the
prescale factor PS8 to 1 and PS1 and 2 to 99999 to do spot++.
- Make sure that the raster is on!! Check that it is at the
correct size.
- After ensuring that the raster is on, put in the hydrogen
target and start taking production data.
- We are really behind in our data goals at this point. Please do
not be shy about calling in the expert on something if it is holding
things up.
Concerns
- Scot Spiegel would like to make an access if we are going to be
down for a while at a time when we can go to controlled access (in the
daytime, 8am-5pm) to get some poly for shielding. If such a time
presents itself, his extension is x5900, his cell is 757-536-5900, and
his pager is 757-584-5900.
- Page Heidi if there is an access in the daytime (6915).
- Q2 quenched yesterday, and Ed asked that we keep a close eye
on it.
- If anyone must reset a magnet, they must make a halog entry
stating which magnet on which arm, and what was wrong with it to require
a reset. The word "tech" must appear in the subject line.
- Be sure that the prompt analysis is done for every run!
- Please log CODA reboots in the halog and shift summary so that it
is easier to track how much time is lost to these reboots.
- The BTA program doesn't subtract any time to speak of for
beam trips. Please manually subtract a full minute for each beam trip in
the beam time accounting.
- Please write a detailed shift summary (including runs taken and any
activities) and submit the BTA for each shift.
- Please make sure that the shift checklist is done once each shift.
- If you have to page or call someone for support, please do so from a
counting house phone or it will be ignored!
- Keep a close eye on the lumi widths (RMS) as we take data.
- Please continue to take runs even if there is no beam. These data
can be used to determine pedestals. Please label these runs as having no
beam.